OpenStudy (anonymous):

Look at the relationship between a and b. a 2 5 8 11 b 7 10 13 16 Which equation below describes the relationship between a and b?

OpenStudy (anonymous):

A. b = 3 × a B. b = 9 - a C. b = 14 ÷ a D. b = 5 + a

OpenStudy (anonymous):

Its D